### v1

@@ 4.4.1 Third-party Domain Knowledge
We emphasized the significance of auditing both the visible external connections and the internal self-regulatory mechanisms of algorithmic systems. NGOs, researchers, journalists, and members of civil society offer domain expertise that is absent in the organizations deploying algorithms [1]. Identifying such contextualities involves a broad range of perspectives about the impacts that such systems might entail. AI experts might not be capable of addressing this issue, and sometimes they are less capable.
According to Li et al. [48], they are not always aware of the way digital technologies affect marginalized communities, and their tools for investigating bias, for example, might not necessarily detect behavior that could not be reasonably predicted before its deployment. Instead, the authors write, “most expert-led auditing methods were developed to detect statistical disparities not, for example, if an algorithm is censuring or harmfully depicting a marginalized community in images or the provision of online services.” This is why we argue that these third parties potentially have specific domain knowledge that include sociotechnical system thinking, as they are have knowledge of human rights impacts in specific domains and are closer to marginalized or affected communities, which brings us to the next aspect.
@@ 4.4.2 Inclusion of Affected Communities
We call this a participatory audit [^75] At this point, it is important to note that the terminology surrounding what we call participatory audits has not yet been consolidated. When affirming the relevance of participatory audits in the context of governments’ accountability,[10] refers to “social audits” in a way that is very similar to what we describe as participatory. According to him, they differ from other forms of audits because they are performed by nonexperts and rely on engagement from citizens and/or civil society organizations. It is a mechanism of oversight, of exerting control on government officials to ensure transparency, responsiveness, and effectiveness by putting citizens in the position of active participants, not mere sources of information [10]., as it consists of a framework that aims to promote the participation of affected stakeholders in different aspects and moments of the design, development, and deployment of AI systems. Audits that consider individual or collective accounts of experiences related to algorithmic bias, harm, or injustice through incident reporting are currently predominantly third-party algorithmic audits [71].
Some authors have cited similar approaches for AI auditing as user-engaged approaches; these would consist of the direct participation of users of AI-based applications in surfacing harmful algorithmic behaviors in activities called “bias bounty” [30], for instance. However, since those affected by AI systems’ supply chains will not always be users of these technologies, we do not consider this to include affected communities [^76] They can also be workers who train AI systems to identify images correctly or in the case of ADMs, where decisions are made about affected people. It is possible that some of these individuals will never use these tools.. In this sense, one can frame participatory audits as a set of different methodologies to include users or affected individuals and groups as a whole in AI auditing processes.
Through methods such as workshops, bias bounty programs, and direct interviews, individuals potentially harmed by these systems can identify problems and help find solutions. NGOs, civil society organizations, and investigative journalists have the opportunity to involve affected communities and users. Of course, companies may have the capability to do the same; however, this appears to be highly improbable with regard to the company’s incentives. Successful community engagement depends on the willingness of the audited organization to genuinely listen, respond, and act on the concerns and recommendations identified during the audit process [1]. Affected communities may trust third-party auditors more than internal audits conducted by the organization itself. This trust can encourage community members to openly share their perspectives, concerns, and experiences during the audit process.
@@ 4.4.3 Information and Power Asymmetries
We have seen in section 4.1.3 that third-party audits can uncover harms despite limited access. To understand what harms are possible, researchers need to access data. This sort of understanding is necessary for new standards-setting. There is an information asymmetry in which data is collected about users and non-users of platforms and algorithmic systems. But at the same, there is no access for researchers to understand the use of the collected d users and non-users of platforms and algorithmic systems is collected data about[89]. Therefore, an audit ecosystem needs leveled access by different auditors who have to publish and share their audit results. Such leveled access can reduce information asymmetry and, ultimately, power asymmetry.
@@ 5.1 Evaluation of the Audit Status Quo in the DSA and the AIA
As noted above, the EU Commission’s AIA proposal relies primarily on internal conformity assessment as a means of assurance. The fact that the focus of the AIA is mainly on self-regulation is questionablebecause there is a lot at stake: high-risk AI systems covered by the AIA can violate numerous legal interests, e. g. people’s health and lives at risk, but also damage property or the environment or violate legal interests such as the protection of privacy. Moreover, implementing the risk-based approach of the AIA requires taking preventive measures that should not solely rely on the assessment of the AI provider, especially for high-risk systems [42].
@@ 5.3 Policy Recommendations
In the delegated acts of the DSA, the inclusion NGOs and investigative journalists as vetted researchers who can perform third-party audits is necessary. Their involvement can ensure social accountability through domain knowledge and special access to accountability by including affected communities. At the same time, journalists have already shown in the past how important they are as an accountability component in financial audits and also in algorithmic audits.
@@
Incorporate provisions for third-party access in the AIA, particularly for ”high-risk” AI systems. These provisions ensure that critical systems undergo rigorous scrutiny and that hard standards can be questioned in the light of new technologies.
@@ 6 Conclusion
We defined audit types – focusing on third-party audits – and analyzed to which extent these audit types and data access are included in the EU regulatory framework.
According to our analysis, the DSA will establish a novel institutional system that involves third-party audits. At the same time, the DSA provides data access for vetted researchers. However, the requirements for independence in the DSA still need to be specified. The European legislator’s AIA proposal relies primarily on internal conformity assessment. At present, access for researchers is not explicitly regulated.
Based on this analysis, we define an AI audit ecosystem that accounts for compliance and oversight and demonstrate the existence of a regulatory gap. We emphasize that third-party audits by research and civil society must be part of that ecosystem. We demand that the AIA include data and model access for certain AI products.
We hope to contribute to the emerging field of audits as a means of accountability. In doing so, we aim to sharpen the focus on affected communities and incorporate past lessons from platform research and financial audits. We recommend establishing an AI audit ecosystem to ensure compliance and oversight. We concur with Raji et al. [70] demands for an ”ecosystem in which third parties can not only survive in their role, but thrive in directly confronting, verifying, and subjecting to scrutiny the performance claims made by corporations, while adequately addressing complaints of harm brought forth by the impacted population.” Our policy proposals point to democratizing accountability, and we hope future AI regulations and the EU directives mentioned here will address them.

### v3

@@ 4.4.1 Third-party Domain Knowledge
We emphasized the significance of auditing both the visible external connections and the internal self-regulatory mechanisms of algorithmic systems. NGOs, researchers, journalists, and members of civil society offer domain expertise that is absent in the organizations deploying algorithms [1]. Identifying such contextualities involves a broad range of perspectives about the impacts that such systems might entail. AI experts might not be capable of addressing this issue.
According to Li et al. [53], AI auditors are not always aware of the way digital technologies affect marginalized communities, and their tools for investigating bias, for example, might not necessarily detect behavior that could not be reasonably predicted before its deployment. Instead, the authors write, “most expert-led auditing methods were developed to detect statistical disparities not, for example, if an algorithm is censuring or harmfully depicting a marginalized community in images or the provision of online services.” This is why we argue that researchers and civil society representatives potentially have specific domain knowledge that include sociotechnical system thinking, as they have knowledge of the human rights impacts that might take place in specific domains exactly for their proximity to marginalized or affected communities. However, including also the communities impacted by the systems, and not those who research or support them, may also prove fundamental for the advancement of systems that duly protect rights.
@@ 4.4.2 Inclusion of Affected Communities
We call this a participatory audit [^75] At this point, it is important to note that the terminology surrounding what we call participatory audits has not yet been consolidated. When affirming the relevance of participatory audits in the context of governments’ accountability,[10] refers to “social audits” in a way that is very similar to what we describe as participatory. According to him, they differ from other forms of audits because they are performed by nonexperts and rely on engagement from citizens and/or civil society organizations. It is a mechanism of oversight, of exerting control on government officials to ensure transparency, responsiveness, and effectiveness by putting citizens in the position of active participants, not mere sources of information [10]., as it consists of a framework that aims to promote the participation of affected stakeholders in different stages of the design, development, and deployment of AI systems. Audits that consider individual or collective accounts of experiences related to algorithmic bias, harm, or injustice through incident reporting are currently predominantly third-party algorithmic audits [77].
Some authors have cited similar approaches for AI auditing as user-engaged approaches; these would consist of the direct participation of users of AI-based applications in surfacing harmful algorithmic behaviors in activities such as those called “bias bounty” [34]. . In this sense, one can frame participatory audits as a set of different methodologies to include users or affected individuals and groups as a whole in AI auditing processes.
Through methods such as workshops, bias bounty programs, and direct interviews, individuals potentially harmed by these systems can identify problems and help find solutions. That is why civil society organizations, researchers and investigative journalists could benefit from involving affected communities and users in their analysis of these technologies. Besides, AI developers and providers should have the capability to do the same, so as to involve impacted groups from the early design of their systems. The same applies to auditors, who can have their assessments of algorithms from having access to other perspectives in order to identify other impacts posed by these technologies. However, despite the potential benefits, there is lack of incentives for such participatory methodologies to be put into place. That is why regulation and incentives should lead developers, deployers and auditors to take this sort of action. Successful community engagement depends on the willingness of the audited organization to genuinely listen, respond, and act on the concerns and recommendations identified during the audit process [1]. Affected communities may trust third-party auditors more than internal audits conducted by the organization itself. This trust can encourage community members to openly share their perspectives, concerns, and experiences during the audit process.
@@ 4.4.3 Information and Power Asymmetries
In section 4.1.3, we learned that third-party audits can uncover harms even with limited access. To understand possible harms, researchers must be able to access data. This understanding is crucial for developing new standards. Currently, there is an imbalance where data is collected about both users and non-users of platforms and algorithmic systems, but there is only limited access for researchers to understand how this data is used [96]. Therefore, an audit ecosystem needs to provide leveled access to different auditors who should publish and share their audit results. This leveled access can help reduce these information and power imbalances.
@@ 4.5 Challenges Of Third-Party Audits
Third-party audits have some drawbacks. These auditors may lack the technical knowledge or capacity to thoroughly research a system. There is also a risk of revealing secret information or company secrets, which can be prevented by including nondisclosure clauses. However, to protect the auditors, safe harbor clauses should be included as well to prevent them from intentional or unknowingly leaking trade secrets. Companies may fear to share their model and data [22]. However, there are best-practices in the case of VLOPs via an API and a vetting process. Further, there are techniques available for adequately protecting or anonymizing data while still providing access and other industries like the finance sector where there are legal ptrotections in the case of data leakage [16].
@@ 5.1 Evaluation of the Audit Status Quo in the DSA and the AIA
As noted above, the EU AIA trilogue version relies primarily on internal conformity assessment as a means of assurance. The fact that the focus of the AIA is mainly on self-regulation is questionable because there is a lot at stake: high-risk AI systems covered by the AIA can violate numerous legal interests, e. g. people’s health and lives at risk, but also damage property or the environment or violate legal interests such as the protection of privacy. Moreover, implementing the risk-based approach of the AIA requires taking preventive measures that should not solely rely on the assessment of the AI provider, especially for high-risk systems [46].
@@ 5.3 Policy Recommendations
In the DSA delegated act on data access, we recommend including NGOs and investigative journalists as vetted researchers who can perform third-party audits. Their involvement can ensure social accountability through domain knowledge and special access to affected communities. Additionally, journalists have demonstrated their importance in financial and algorithmic audits as an accountability component in the past.
@@
Incorporate provisions for third-party access in the AIA, particularly for “high-risk” AI systems. These provisions ensure that critical systems undergo rigorous scrutiny and that hard standards can be questioned in the light of new technologies.
@@ 6 Conclusion
We defined audit types – focusing on third-party audits – and analyzed to which extent they are included in the EU regulatory framework, with a focus on the DSA and the AIA.
According to our analysis, the DSA will establish a novel institutional system that involves third-party audits, provides data access for vetted researchers as well as access to openly available data for non-vetted researchers. However, the requirements for independence of auditors in the DSA still need to be specified. The AIA, on its turn, relies primarily on internal conformity assessment with external assessments that can be mandated by the regulator. At present, access for researchers or independent auditors is not explicitly intended.
Based on this analysis, we define an AI audit ecosystem that accounts for compliance and oversight and demonstrate the existence of a regulatory gap. We emphasize that third-party audits by independent organisations, researchers and civil society, as well as through inspection by regulators, must be part of that ecosystem. The AIA, in this sense, should have included data and model access for vetted researchers and civil society organisations to assess “high-risk” AI products, as is partially the case of the DSA’s provisions allowing for vetted researchers — although not civil society organisations — to have access to sensitive information on the AI systems.
We hope to contribute to the emerging field of audits as a means of accountability. In doing so, we aim to sharpen the focus on affected communities and incorporate past lessons from platform research and financial audits. We recommend establishing a diverse AI audit ecosystem to ensure compliance and oversight. We concur with Raji et al. [76]’s demands for an “ecosystem in which third parties can not only survive in their role, but thrive in directly confronting, verifying, and subjecting to scrutiny the performance claims made by corporations, while adequately addressing complaints of harm brought forth by the impacted population.” Our policy proposals point to democratizing accountability, and we hope future AI regulations and the EU directives mentioned here will address them.